OpenStudy (anonymous):

Find the inverse function of f(x)=x-4/7

OpenStudy (mathmale):

Have you found inverse functions before? if so, what steps did you go through to find the inverse of a given function?

OpenStudy (anonymous):

f^−1(x)=4/7+x

OpenStudy (anonymous):

inverse function is the opposite of the function

OpenStudy (whpalmer4):

@sensuelle1985 uh, no, that isn't correct.

OpenStudy (anonymous):

can you explain why please? @whpalmer4

OpenStudy (whpalmer4):

I was hoping that the original poster would answer my colleague's question... The way I find inverse functions is by swapping the variables and solving. As an example, I'll do a problem I just did elsewhere on OS: \[y = f(x) = \frac{3x-4}{5}\]We swap the variables: \[x = \frac{3y-4}{5}\]Now solve for \(y\):\[5x=3y-4\]\[5x+4=3y\]\[y=f^{-1}(x) = \frac{5x+4}{3}\]

OpenStudy (whpalmer4):

One of the nifty things about finding inverse functions is that they are the same as the original function, except reflected across the line \(y = x\):

OpenStudy (whpalmer4):

In that graph, the purple and blue lines are the function and inverse (I don't recall which is which), and the olive line is the line of reflection

OpenStudy (anonymous):

Okay got it, thanks for the clarification but is my answer right though?

OpenStudy (whpalmer4):

No, it's not. Assuming you mean \[f^{-1}(x) = \frac{4}{7} + x\]that's just \(y=x\) shifted up by \(\large \frac{4}7\), not a reflection of \(\large y = \frac{x-4}7\), which is undoubtedly what the poster intended. Yes, yours is correct if you interpret what they wrote "correctly" but the usual circumstances here are students don't realize that \[x+4/y\] does not mean \[\frac{x+4}7\]